A three-year-old machine that has a cost of $63,000, an estimated residual value of $5,800, and an estimated useful life of five years. The company uses double-declining-balance depreciation.

Calculate the net book value at the end of each year.

Net book value:

First year

Second year

Third year
Answer #1

Annual depreciation rate as per straight line method=100/5=20%/year

Hence depreciation as per double decline balance=2*Annual depreciation rate as per straight line method*Beginning value of each period

Year Beginning value Depreciation Net book value at end of year
1 63000 (2*20%*63000)=$25200 (63000-25200)=$37800
2 37800 (2*20%*37800)=$15120 (37800-15120)=$22680
3 22680 (2*20%*22680)=$9072 (22680-9072)=$13608.